OpenStudy (anonymous):

giving medal Given the formula a-b+1=c, solve for b. a.-b=c-a-1 c. b=-c-a-1 b. b=-c+a+1 d. b=-c=a-1

OpenStudy (anonymous):



OpenStudy (anonymous):

no I change it to C.!

OpenStudy (anonymous):

you got anymore questions??

OpenStudy (anonymous):

@Torrayani<3 hpw did u get that explain pl

OpenStudy (anonymous):

because if you notice in this problem (a-b+1=c), a=-a -b=b 1=-1, and c=-c and answer c. b=-c-a-1 was the only one that got that whole situation right @shkrina